OpenStudy (anonymous):

F = 27.2•m ------- Solve t

OpenStudy (diyadiya):

What do you have to solve?

OpenStudy (anonymous):

the problem

OpenStudy (diyadiya):

What is F ,m & t ? do you have to solve it for a certain variable?

OpenStudy (anonymous):

no just the problem

OpenStudy (diyadiya):

Is there anything else given?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh wait yes solve for M

OpenStudy (diyadiya):

Well, yeah that was what i was asking for xD

OpenStudy (diyadiya):

Step1: Multiply both sides by t Step 2: Divide both sides by 27.2

OpenStudy (anonymous):

uhhhhhhhh......?

OpenStudy (diyadiya):

First multiply both sides by t what do you get now?

OpenStudy (anonymous):

27.2t?

OpenStudy (diyadiya):

Both sides \[F \times t = \frac{ 27.2 \times m }{t} \times t\]

OpenStudy (diyadiya):

now what happens? does anything cancel out?

OpenStudy (anonymous):

umm not shure :(

OpenStudy (diyadiya):

Ok i can show you this step but you'll have to show me the next step.

OpenStudy (diyadiya):

since t/t = 1\[F \times t = \frac{ 27.2 \times m }{\cancel{t}} \times \cancel{t}\]\[F \times t = 27.2 \times m \] basically what we're trying to do is to isolate m so now we have 27.2 with m on the right side what can you do with that? Btw did you understand till this part?

OpenStudy (anonymous):

kinda

OpenStudy (diyadiya):

Ok now? :D

OpenStudy (diyadiya):

Now divide both sides by 27.2 :)